Fermi National Laboratory
 

Accelerator Update for April 24, 2009 - April 27, 2009

 

The Integrated Luminosity 4/20/09 to 4/27/09 was 47.42 pb-1.

Friday April 24

The day shift began with Operations monitoring the Tevatron (TeV) store 6993 (75.82E30) and the Recycler stash (349.3E10), with the Antiproton Source stacking (22.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Operations terminated store 6993 at 8:30 AM.

Operations prepared the Transfer Hall, TeV sectors A-E, and F-sector for access to address vacuum problems at sector F23.

MI-60 suffered a vacuum burst at 9:00 AM and was recovered shortly thereafter.

Operations had established beam to the Pbar, NuMI, and MiniBooNE by 12:30 PM.

Operations had put the TeV into HEP shot setup by 3 PM.

Operations established store 6997 at 4:30 PM with an initial luminosity of 343.4E30.

A Cold Box at CHL failed at 5 PM. The TeV sector A4 cryo system went unstable leading to a quench and loss of store 6997.

Operations delivered beam to the Pbar, NuMI, and MiniBooNE until a Linac RF station (LRF2) failed at 10:30 PM.

CHL personnel made the necessary repairs and restarted the cold box by 11 PM.

MIRF personnel were called in to replace the bias power supply for a MI RF station (MIRF6).

Antiprotons transfers to the Recycler were completed when necessary.

Saturday April 25

The midnight shift began with the TeV recovering from a cold box failure, with Operations monitoring the stash (165.5E10) and the stack (11.7E10), with Linac down due to LRF2 problems, and with NuMI and MiniBooNE off due to the Linac problems.

The TeV sector E0 crystal collimator was removed and replaced with a spool piece.

The MI vertical sextupole power supply developed an internal problem; EE Support personnel were called in and switched over to the backup power supply.

A LCW supply line to the I- Source water resistorfailed at a compression fitting. Linac personnel were called in to make repairs.

A lightning glitch tripped the MI ramp off on two occasions. 50E10 of the Recycler stash was lost in reestablishing and re-learning the MI ramp.

The Electron Cooling doublet magnet at the bottom of the Pelletron tank required a back flush of its LCW cooling line. Waterworks personnel were called in to perform the task.

Linac techs completed the LRF2 repairs and the station was tuned up. Beam operations began in earnest at 8:30 PM.

The TeV ring cooldown progressed. Cryo system engine techs were called in to address wet engine problems at TeV sector F4.

DZero accessed their collider hall.

Beam was delivered to the Pbar Source, NuMI, MiniBooNE, and the Meson MTest when available.

Antiprotons transfers to the Recycler were completed when appropriate.

Sunday April 26

The midnight shift began with TeV cooldown in progress, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The TeV cryo system had recovered by noon, and machine turn-on and checkout began. Before getting too far into machine turn-on, operators accessed the TeV A-E sectors to check out the status of the C3 sump. Upon resuming the TeV checkout, a hipot revealed a ground fault between sectors F2 and A2. Operations consulted TeV personnel and a ground fault hunt ensued. Nothing obvious was found; operators resumed the TeV check out and the turn-on resumed after the fault disappeared.

Antiprotons transfers to the Recycler were completed when appropriate, leading to a New Recycler stash record of 502.85E10.

Operations put the TeV through a dry squeeze and orbit check at 10:30 PM.

Operations had put the TeV into HEP shot set-up by 10:45 PM.

Beam was delivered to the Pbar, NuMI, MiniBooNE, and MTest when available.

Monday April 27

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established store 7000 at 12:15 AM with an initial luminosity of 304E30.

The stash of ~315E10 was lost during a C453 controls card swap.

Beam was delivered to the Pbar, NuMI, MiniBooNE, and MTest when available.

Antiproton transfers to the Recycler were completed when appropriate.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/28/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory